Advanced Accounting Ch 1 Q 6

Bradford, Inc., acquires 30 percent of the outstanding voting shares of Simmons, Inc., on January 1, Year 10, for $312,000, which gives Bradford the ability to significantly influence Simmons.  Simmons has a net book value of $800,000 at January 1, Year 10.  Simmons’s asset and liability accounts showed carrying amounts considered equal to fair values, except for a copyright whose value accounted for Bradford’s excess cost over book value in its 30 percent purchase.  The copyright had a remaining life of 16 years at January 1, Year 10.  No goodwill resulted from Bradford share purchase. 

    Simmons reported net income of $180,000 in Year 10 and $230,000 of net income during Year 11. Dividends of $70,000 and $80,000 are declared and paid in Years 10 and 11, respectively.  Bradford uses the equity method. 


a. On its Year 11 comparative income statements, how much income would Bradford report for Year 10 and Year 11 in connection with the company’s investment in Simmons? 


b. If Bradford sells its entire investment in Simmons on January 1, Year 12 for $400,000 cash, what is the impact on Bradford’s income? 


c. Assume that Bradford sells inventory to Simmons during Years 10 and 11 as follows: 

   ____________________________________________________________________________

                                  Cost to                        Price to                       Year-End Balance 

        Year               Bradford                      Simmons                       (at transfer price) 

          10                  $30,000                         $50,000                    $20,000 (sold in following year)

          11                    33,000                           60,000                      40,000 (sold in following year)

              

           What amount of equity income should Bradford recognize for Year 11? 
Answer #1

Step 1: Find yearly amortization. 

Yearly amortization for the copyright = purchase price - net book value / remaining life of the copyright


[312,000 - (800,000 x 30%)] /  16 years = 4,500 yearly excess amortization


Step 2:

a)


Find equity income for both years.

Year 10 net income = (180,000 x 30%) - 4,500 amortization = 49,500

Year 11 net income = (230,000 x 30%) -  4,500 amortization = 64,500


Step 3: 

Find dividends for both years

Year 10 dividend = 70,000 x 30% = 21,000

Year 11 dividend =  80,000 x 30% = 24,000


Step 4: Find deferred gross profit for Year 10 and Year 11

Deferred gross profit = ending inventory x gross profit % x ownership %

Year 10  gross profit % = (50,000 price to simmons - 30,000 cost to bradford) / 50,000 = 40% gross profit

= 20,000 x 40% x 30% = 2,400


Year 11 gross profit % = (60,000 price to simmons - 33,000 cost to bradford) / 60,000 = 45%

40,000 x 45% x 30% = 5,400


Step 4: Find gain on sale

Book value = 

[312,000 original book value + 54,0000 net income Y10 - 21,000 dividend Y10 - 4,500 excess amortization Y10 - 2400 deferred gross profit Y10] + [69,000 net income Y11 + 2400 sold Y11 - 24,000 dividends Y11 - 4,500 excess amortization Y11] = 381,000 book value


b) Gain on sale = 400,000 sales price - 381,000 book value = 19,000


c) Equity income = Net income for year 11 + deferred gross profit year 10 - deferred gross profit year 11. We add the deferred gross profit for year 10 because we sold it in year 11, and we subtract deferred gross profit for year 11 because we have not sold it yet.


64,500 + 2,400 - 5,400 = 61,500 




answered by: anonymous